OpenStudy (anonymous):

What is the sum of 2 1/4 and 6 2/5 ?

OpenStudy (anonymous):

\(\Large 2\frac{1}{4}=\frac{2\times4+1}{4}=\frac{9}{4}\) And \(\Large 6\frac{2}{5}=\frac{6\times5+2}{5}=\frac{32}{5}\) Thus, \(\Large 2\frac{1}{4}+6\frac{2}{5}=\frac{9}{4}+\frac{32}{5}=\frac{9\times5+32\times4}{4\times5}=\frac{173}{20}\)